When NMOS and PMOS are used as pass transistors, briefly explain why the NMOS transistor passes a weak digital “1” and the PMOS transistor passes a weak digital “0”.
Answer #1

NMOS ON when logic HIGH (5 volts) is given at GATE

NMOS has negative threshold voltage

Vgs>=Vth , Vs will be passed at Vd

consider for example Vth is 1

Vs= 0, Vgs=5-0= 5 ,Vgs>Vth, ON, Vd=0      (NMOS passes strong 0)

Vs= 3, Vgs=5-3= 2, Vgs>Vth, ON, Vd=3

Vs=4 , Vgs=5 -4=1 ,Vgs=Vth, barely ON , Vd=4

Vs= 5, Vgs=5-5= 0 ,Vgs<Vth, OFF , Vd=4(previous value passed) (NMOS passes weak 1)

PMOS ON when logic LOW (0 Volt) is given at GATE

PMOS has negative threshold voltage

Vgs<=Vth , Vs will be passed at Vd

consider for example Vth is -1

Vs= 5, Vgs=0-5= -5 ,Vgs<Vth, ON, Vd=5 (PMOS passes strong 1)

Vs= 2, Vgs=0-2= -2 ,Vgs<Vth, ON, Vd=2

Vs= 1, Vgs=0-1= -1 ,Vgs=Vth, barely ON , Vd=1

Vs= 0, Vgs=0-0= 0 ,Vgs>Vth, OFF , Vd=1(previous value passed) (PMOS passes weak 0)